# On the Onset of Coherent Phonon Motion in Peierls-Distorted Antimony by Attosecond Transient Absorption

### v2

@@ Appendix B DFT Calculations
Band structure and (projected) density of states were calculated using the Quantum Espresso software package. An ultrasoft pseuopotential with Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional from the GBRV library was used 52. Convergence of the self-consistent field calculations was found at a wavefunction cutoff of 40 Ry and a grid of 203 points in k-space in accordance with the standard solid-state pseudopotentials (SSSP) library 53. The lattice was relaxed using the variable cell subroutine, leading to lattice constants of and , and a Peierls parameter , in good agreement with crystallographic parameters.
@@ Appendix C Experimental Setup
To control the temporal delay between XUV and NIR pulses, a retroreflector is mounted atop two motorized stages in the pump arm, a long travel stepper-motor controlled stage and a short travel piezo-electric stage. While the stepper-motor controlled stage allows to scan picosecond ranges, the piezo-electric stage allows finer delay control on the sub-femtosecond scale. To avoid artifacts from variances in the XUV spectrum, static XUV spectra are rapidly taken sequentially for the sample and blank membranes. Similarly for pump-probe spectra, a mechanical shutter in the pump arm allows to rapidly take sequential spectra with and without the pump beam present. The spectra taken without the pump beam are used to calculate a noise correlation matrix from regions of the spectrum where no change of absorbance is observed, and this is used to suppress correlated source noise in the observed transient spectra 54 (see below).
@@ Appendix D Time-Zero and Instrument Response Function Fitting
To calibrate the XUV-NIR zero delay, attosecond transient absorption spectra of AC Stark shifting of the Ne 3s4p autoionization resonance are recorded periodically during the delay-dependent scans. The light-induced phase signal of the autoionization line is analyzed through SVD and the zero-delay is obtained through a non-linear least-squares fitting of a Gaussian instrument response function (IRF) convoluted with the exponential decay of the autoionizing resonance 46. In difference to the pump-probe spectra recorded for Sb, the Ne spectra where recorded only with the pump present. To suppress the noise originating in a change of the harmonic spectrum, the spectra are high-pass filtered, since the relevant autoionizing lines of Ne are much narrower than the typical harmonic noise 44. Fig. 5a shows the (averaged) counts of a typical Ne scan (blue) with the low-passed portion of the spectrum (orange) representing the typical HHG spectrum during the experiment. Fig. 5b shows the attosecond transient absorption spectrum of Ne after high-pass filtering as a change in absorbance, i.e. after subtracting the static absorption. The temporal component of one exemplary scan is shown in Fig. 5c with a fitted exponentially modified Gaussian distribution, yielding an IRF of fs.
@@ E.1 Data loading, Edge-Referencing and Heat Subtraction
Fig. 6(a) shows the average change of absorption per delay before edge-referencing and Fig. 6(b) afterwards. Before zero-delay, a small residual signal can be seen. The signal, which is interpreted as residual heat lingering between consecutive pump-pulses, is averaged for the first 20 delay points of the scan ( fs) and subtracted from all transient absorption spectra. Fig. 6(c) shows the subtracted signal. It is comparable to the signal at the end of the long delay scan ( ps), albeit roughly 3x smaller in intensity, supporting our interpretation as residual heat. Previous investigation of this laser induced heating signal has shown that the removal of the feature does not significantly affect the measured dynamics beyond the effect of an elevated equilibrium temperature 55, 13. For further analysis, the spectral axis is transformed from wavelength (nm) to photon energy (eV) by linear interpolation.
@@ E.2 Spin-Orbit Deconvolution
To recover the spectral response of a single core-subshell, we perform spin-orbit deconvolution by Fourier transform 13: It is assumed that the measured spectra follow a convolution of each spin-orbit component of the core-level, such that the measured spectrum :
@@ Appendix G Extended Two Temperature Model
where is the carrier heat capacity, the electron-phonon coupling term with the lattice temperature . The last term accounts for the back-coupling effect due to the change of potential energy caused by the phonon motion per unit volume 30.
@@ Appendix I Amplitude of Coherent Phonon Mode with Carrier-Relaxation
We will start with the general solution of the differential equation (cf. Eqn. (1,2)). Again, following the work by O’Mahony et al.11, we solve using the initial conditions and , which gives us:
@@
This result is very similar to that reported in Ref. 11, i.e. it represents a lowering of the phonon amplitude due to the time-dependent force and is dependent on the scaling parameter . This implies that the results that were obtained in Ref. 11 assuming that a decrease in force on the A1g-mode during carrier relaxation could be compatible with the assumed increase in force in this work.
@@
As an alternative approach to evaluate the phase of the phonon motion, a Fourier analysis using a Takeda-algorithm is performed 56, 57. The frequency dependent amplitude for each photon energy is calculated by a discrete Fourier transform. Due to the strong exponential rise- and decay- dynamics of the carrier relaxation, a strong background tail extending out from the DC-component is observed, limiting the contrast at the phonon peak. The contrast can be severely enhanced and the background eliminated by subtracting the purely thermal model obtained through the component analysis from the measured data. The resulting Fourier spectogram is shown in Fig. 10a. The phonon motion is sharply peaked around THz, in good agreement with the result of the fit (see Fig. 10a).

### v3

@@ Appendix B DFT Calculations
Band structure and (projected) density of states were calculated using the Quantum Espresso software package. An ultrasoft pseuopotential with Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional from the GBRV library was used [56]. Convergence of the self-consistent field calculations was found at a wavefunction cutoff of 40 Ry and a grid of 203 points in k-space in accordance with the standard solid-state pseudopotentials (SSSP) library [57]. The lattice was relaxed using the variable cell subroutine, leading to lattice constants of and , and a Peierls parameter , in good agreement with crystallographic parameters.
@@ Appendix C Experimental Setup
To control the temporal delay between XUV and NIR pulses, a retroreflector is mounted atop two motorized stages in the pump arm, a long travel stepper-motor controlled stage and a short travel piezo-electric stage. While the stepper-motor controlled stage allows to scan picosecond ranges, the piezo-electric stage allows finer delay control on the sub-femtosecond scale. To avoid artifacts from variances in the XUV spectrum, static XUV spectra are rapidly taken sequentially for the sample and blank membranes. Similarly for pump-probe spectra, a mechanical shutter in the pump arm allows to rapidly take sequential spectra with and without the pump beam present. The spectra taken without the pump beam are used to calculate a noise correlation matrix from regions of the spectrum where no change of absorbance is observed, and this is used to suppress correlated source noise in the observed transient spectra [51] (see below).
@@ Appendix D Time-Zero and Instrument Response Function Fitting
To calibrate the XUV-NIR zero delay, attosecond transient absorption spectra of AC Stark shifting of the Ne 3s4p autoionization resonance are recorded periodically during the delay-dependent scans. The light-induced phase signal of the autoionization line is analyzed through SVD and the zero-delay is obtained through a non-linear least-squares fitting of a Gaussian instrument response function (IRF) convoluted with the exponential decay of the autoionizing resonance [47]. In difference to the pump-probe spectra recorded for Sb, the Ne spectra where recorded only with the pump present. To suppress the noise originating in a change of the harmonic spectrum, the spectra are high-pass filtered, since the relevant autoionizing lines of Ne are much narrower than the typical harmonic noise [45]. Fig. 6a shows the (averaged) counts of a typical Ne scan (blue) with the low-passed portion of the spectrum (orange) representing the typical HHG spectrum during the experiment. Fig. 6b shows the attosecond transient absorption spectrum of Ne after high-pass filtering as a change in absorbance, i.e. after subtracting the static absorption. The temporal component of one exemplary scan is shown in Fig. 6c with a fitted exponentially modified Gaussian distribution, yielding an IRF of fs.
@@ E.1 Data loading, Edge-Referencing and Heat Subtraction
Fig. 7(a) shows the average change of absorption per delay before edge-referencing and Fig. 7(b) afterwards. Before zero-delay, a small residual signal can be seen. The signal, which is interpreted as residual heat lingering between consecutive pump-pulses, is averaged for the first 20 delay points of the scan ( fs) and subtracted from all transient absorption spectra. Fig. 7(c) shows the subtracted signal. It is comparable to the signal at the end of the long delay scan ( ps), albeit roughly 3x smaller in intensity, supporting our interpretation as residual heat. Previous investigation of this laser induced heating signal has shown that the removal of the feature does not significantly affect the measured dynamics beyond the effect of an elevated equilibrium temperature [58, 13]. For further analysis, the spectral axis is transformed from wavelength (nm) to photon energy (eV) by linear interpolation.
@@ E.2 Spin-Orbit Deconvolution
To recover the spectral response of a single core-subshell, we perform spin-orbit deconvolution by Fourier transform [13]: It is assumed that the measured spectra follow a convolution of each spin-orbit component of the core-level, such that the measured spectrum :
@@ Appendix G Extended Two Temperature Model
where is the carrier heat capacity, the electron-phonon coupling term with the lattice temperature . The last term accounts for the back-coupling effect due to the change of potential energy caused by the phonon motion per unit volume [31].
@@ Appendix I Amplitude of Coherent Phonon Mode with Carrier-Relaxation
We will start with the general solution of the differential equation (cf. Eqn. (1,2)). Again, following the work by O’Mahony et al.[11], we solve using the initial conditions and , which gives us:
@@
This result is very similar to that reported in Ref. [11], i.e. it represents a lowering of the phonon amplitude due to the time-dependent force and is dependent on the scaling parameter . This implies that the results that were obtained in Ref. [11] assuming that a decrease in force on the A1g-mode during carrier relaxation could be compatible with the assumed increase in force in this work.
@@
As an alternative approach to evaluate the phase of the phonon motion, a Fourier analysis using a Takeda-algorithm is performed [59, 60]. The frequency dependent amplitude for each photon energy is calculated by a discrete Fourier transform. Due to the strong exponential rise- and decay- dynamics of the carrier relaxation, a strong background tail extending out from the DC-component is observed, limiting the contrast at the phonon peak. The contrast can be severely enhanced and the background eliminated by subtracting the purely thermal model obtained through the component analysis from the measured data. The resulting Fourier spectogram is shown in Fig. 11a. The phonon motion is sharply peaked around THz, in good agreement with the result of the fit (see Fig. 11a).